In a certain Algebra 2 class of 28 students, 7 of them play basketball and 5 of them play baseball. There are 18 students who play neither sport. What is the probability that a student chosen randomly from the class plays both basketball and baseball?



1
Expert's answer
2021-11-04T12:27:57-0400

Number of students = 28

7 of them play basketball and 5 of them play baseball. There are 18 students who play neither sport.

Number of students playing = 28 – 18 = 10

Number of students playing both games = 7 + 5 – 10 = 2

Probability that a student chosen randomly from the class plays basketball and baseball:

"P = \frac{2}{28} = 0.0714"

Answer: 0.0714